OpenStudy (anonymous):

Help! I'm working on this quadratic equation: 5x^2 = -30x - 65 I've got most of it understood, only I've reached a part in the equation where I need some help. Here: 30 +/- 2i sqrt(100) / 10 I'm not sure how to progress!

OpenStudy (immanuelv):

just simplify

OpenStudy (anonymous):

Simplify 100?

OpenStudy (immanuelv):

sqrt of 100 = 10

OpenStudy (anonymous):

Oh-! Okay, thank you! would I add that to 2i then? or.. I think 2i cancels out?

OpenStudy (immanuelv):

hold on i got \[x = -30 \pm \frac{ \sqrt{-1210} }{10 }\]

OpenStudy (immanuelv):

it will be 11i sqrt 10

OpenStudy (anonymous):

How did you get that??

OpenStudy (immanuelv):

quadratic formula

OpenStudy (anonymous):

\[5x^2 = -30x - 65\\ 5x^2+30x+65=0\\ x^2+6x+13=0\] would be a good start

OpenStudy (anonymous):

then since \(6\) is even, it would be easiest to complete the square an ignore the formula \[x^2+6x+13=0\\ x^2+6x=-13\\ (x+3)^2=-13+9=-4\]

OpenStudy (anonymous):

you get \[x+3=\pm\sqrt{-4}=\pm2i\]

OpenStudy (anonymous):

making \[x=-3\pm 2i\]

OpenStudy (anonymous):

which is exactly what you had got in your first answer, if you divide each term by \(10\)

OpenStudy (anonymous):

Oh thank you, that makes more sense!!